OpenStudy (anonymous):

a rectangle who's perimeter is 167 feet, has a length that is 6 feet longer than its width. Find the length and width of the rectangle. im so confused please help

OpenStudy (anonymous):

Let's say the rectangle has side lengths a and b. We know a+6=b, and 2(a)+2(b)=167, or subbing in b=a+6 we get 2(a)=2(a+6)=167. So you can solve for a now and then since a+6=b, you can get b. Notice that a=width, and b=length.

OpenStudy (anonymous):

Ohh thank you sooo much!!!

OpenStudy (anonymous):

*that's 2(a)+2(a+6)=167. I hit the = sign instead of +

OpenStudy (anonymous):

yea i got it thanks!